OpenStudy (anonymous):

Solve this linear system by using elimination 4x+5y=14 2x-3y=-2 Please help and explain in full steps

OpenStudy (anonymous):

1. Multiply one of the equations to ensure that both equations have at least one common factor. E.g. 4x+5y=14 4x-6y=-4 2. Top eqn. - bottom eqn. E.g. -y=10 so y=-10 3. Substitute result into one of the original eqns. E.g. 4x+5(-10)=14 = 4x-50=14 = 4x=64 = x=26. So y=-10 & x=26

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

No problem. I enjoyed it :P

OpenStudy (anonymous):

on my awnser sheet it says x=16/11 and y=18/11

OpenStudy (anonymous):

...let me check my working lol

OpenStudy (anonymous):

Ah... 1. Multiply one of the equations to ensure that both equations have at least one common factor. E.g. 4x+5y=14 4x-6y=-4 2. Top eqn. - bottom eqn. E.g. 11y=18 so y=18/11 3. Substitute result into one of the original eqns. E.g. 4x+5(18/11)=14 = 4x+90/11=14 = 4x=64/11 = x=16/11. So y=18/11 & x=16/11. Yeah, that's better. Made a simple error at part 2 the 1st time.